OpenStudy (anonymous):

The length of a rectangle is 3 yards less than twice the width, and the area of the rectangle is 27 yards^2 . Find the dimensions of the rectangle.

OpenStudy (abmon98):

width=x length=2x-3 x(2x-3)=27 2x^2-3x-27 x=9/2 and x=-3 width=9/2 length=2(9/2)-3=6